OpenStudy (hitaro9):

Find all the values of n such that phi(phi(n)) = 1. How does this relate to primitive roots and what does this tell us?

OpenStudy (rational):

suppose \(n\) has a primitive root, then what can you tell about the number of primitive roots of \(n\) ?

OpenStudy (hitaro9):

I'm not sure? It has at least 1?

OpenStudy (rational):

it should be easy, just wok it using the result from previous problem

OpenStudy (rational):

let \(r\) be a primitive root of \(n\), then all the positive integers less than \(n\) that are relatively prime to \(n\) are given by \[r^1, r^2, r^3\ldots, r^{\phi(n)}\] yes ?

OpenStudy (hitaro9):

Right that's just the def. of primitive root.

OpenStudy (rational):

remember that a primitive root of \(n\) has order \(\phi(n)\)

OpenStudy (rational):

how many integers in above list have order \(\phi(n)\) ?

OpenStudy (rational):

you want to work out : \[\dfrac{\phi(n)}{\gcd(\phi(n), a)} = \phi(n)\]

OpenStudy (rational):

which is same as \(\gcd(\phi(n),a) = 1\) how many positive integers less than \(\phi(n)\) satisfy above relation ?

OpenStudy (rational):

in other words, how many positive integers less than \(\phi(n)\) are relatively prime to \(\phi(n)\) ?

OpenStudy (hitaro9):

Okay, I'm following what you're saying, but not how it relates to this problem.

OpenStudy (rational):

There are exactly \(\phi(\phi(n))\) positive integers less than \(\phi(n)\) that are relatively prime to \(\phi(n)\), yes ?

OpenStudy (hitaro9):

Right

OpenStudy (rational):

that means there are exactly \(\phi(\phi(n))\) primitive roots for any given integer \(n\), if it has a primitive root.

OpenStudy (hitaro9):

Oh okay. I think I see

OpenStudy (rational):

finding all the values of \(n\) such that \(\phi(\phi(n))=1\) is same as finding all the integers having exactly one and only one primitive root.

OpenStudy (hitaro9):

Right that makes sense. Thank you again for the help.

OpenStudy (hitaro9):

I have 1 other problem while you're still here if you wouldn't mind taking a look?

OpenStudy (rational):

I'll try, post..

OpenStudy (hitaro9):

It's pretty similar to a problem in the book, and I followed the steps pretty closely, but when I plug the problem in to wolfram alpha it returns a different answer

OpenStudy (hitaro9):

So we're given this indice table 16,14,1,12,5,15,11,10,2,3,7,13,4,9,6,8

OpenStudy (hitaro9):

for 1-16 mod 17

OpenStudy (hitaro9):

and are asked to find all solutions to 9x^8 = 8 mod 17

OpenStudy (rational):

is the primitive root 3 ?

OpenStudy (hitaro9):

Yes

OpenStudy (hitaro9):

Right

OpenStudy (rational):

\[9x^8 \equiv 8 \pmod {17}\] take discrete logarithm both sides

OpenStudy (hitaro9):

So I took the ind3 on both sides and got ind3(9x^8) = 0 mod 16 14 + 8ind3x = 0 mod 16

OpenStudy (hitaro9):

But that doesn't yield what wolfram alhpa is giving for the original problem

OpenStudy (rational):

how did u get 14 ?

OpenStudy (hitaro9):

(and it has the correct answers for the book problem similar to this one)

OpenStudy (hitaro9):

I used a theorem and broke ind3 (9x^8) in to ind3(9) + ind3(x^8) ind3(9) =14

OpenStudy (rational):

index of 9 is 2 right ? because 9 = 3^2

OpenStudy (hitaro9):

Oh wait. I think I was just reading the table backwards

OpenStudy (rational):

14 is the index of 2 : 2 = 3^14

OpenStudy (hitaro9):

Wow I feel really dumb. Okay that was probably what the issue was, thank you.

OpenStudy (rational):

remember index = logarithm = exponent

OpenStudy (rational):

2 is the index of 9 relative to primitive root 3 because \[\large 9 \equiv 3^{\color{blue}{2}}\pmod{17}\]

OpenStudy (hitaro9):

Right. I'm not quite sure what I was thinking there, thank you.

OpenStudy (hitaro9):

Okay yeah, the problem is looking a lot nicer now that I'm going through and correcting it. Thanks a ton for all your help today.

OpenStudy (rational):

np:)